2023 June

From EOVSA Wiki
Jump to navigation Jump to search

Back to Observing Log

June 07

16:00 UT Calibrations did not run for June 6. The LNA voltages on the 27m were all zero. I rebooted the brick, initialized the LNA and homed the receiver.

June 08

17:30 UT Frequency Tuning had locked up. Trying to reboot did not work. The ACC VI could not be started. The Windows computer was rebooted, as was the ACC before observations were resumed. About 1 hour of observing time was lost.

June 13

23:00 UT The 27-m control system started acting up (immediate symptop was 0.0 LNA readings on the CryoRX display). The problem was traced to a bad TCP communication with the feanta computer, although all components could be pinged successfully. No evening reference calibration was possible.

June 14

12:24 UT This morning I cycled power to the feanta computer and reset the brick controller. Everything is now working again, but not until 12:23 UT or so, which means this morning's reference calibration will be no good.

15:06 UT The data was stuck recording as phasecal even when the array was pointing at the Sun. Restarting the schedule rectified the problem. About 90 minutes of data was lost.

June 15

19:30 UT Dynamic spectrum looked bad. ADC_plot was restarted but made no improvement.

June 16

19:50 UT The 27m antenna would not move. There was a hydraulic fluid alarm and cable tension alarm. Mark Hodges found several blown fuses. On starting the hydraulic pump, it made a clunking sound. The antenna stowed and no calibrations will be done until it can be inspected on the 20th.

June 20

16:53 UT After inspection of the 27m Hydraulic pump, it was determined that it had failed. A replacement will be ordered. Until the pump is replaced, no calibrations can be performed.

June 23

17:24 UT The ACC locked up. It required a restart.

June 26

19:10 UT The DPP was unreachable from any other machine, and no data was being recorded. It required a complete power down and restart.